Five years ago, the age of Samyak was four times the sum of ages of his two children. If twice the present age of Samyak is thirteen years less than five times the sum of the present ages of his two children, then seven years hence the age of Samyak will be
 

Grade:10

1 Answers

Harshit Singh
3 years ago
Dear Student

let the present age of samyak is x
sum of present ages of children = y
five year ago, age of samyak= x –5
five year ago,sum of ages of children =y-10 since there are two children
given,
x –5= 4(y – 10)
x – 5=4y –40
x –4y= –35............(i)
also it is given that
2x=5y –13
2x –5y = –13......(ii)
solving equation i and ii
x= 41 years
y = 19 years

Hence, age of samyak will be 41 years

Thanks